·       An analyst forecasts to have expected dividends of $.75 one year from now for Turtle Company. The dividends are expected to grow at 12% per year thereafter until the fourth year. After the fourth year the dividends will grow at a constant rate of 2% per year. Equity Cost of Capital is 5% Re. Calculate the current stock price P0 using the Dividend Discount Model modified formula with (DivN+1/Re-G) on the end resulting in Pn Stock Price.
Answer #1

Stock price = (D1/(1+R)^1))+((D2/(1+R)^2))+((D3/(1+R)^3))+((D4/(1+R)^4))+(D4*(1+g)/(r-g))/(1+R)^4))

stock price = (D1/(1+R)^1))+((D1*(1+g))/(1+R)^2))+((D1*(1+g)^2)/(1+R)^3))+((D1*(1+g)^3)/(1+R)^4))+(((D1*(1+g)^3)*(1+g))/(r-g))/(1+R)^4))

=((0.75/(1.05))+((0.75*(1.12))/(1.05^2))+((0.75*(1.12^2))/(1.05^3))+((0.75*(1.12^3))/(1.05^4))+((((0.75*(1.12^3)*(1.02))/(0.05-0.02)))/(1.05^4)) = 32.63
Answer #2

SOLUTION :


(Values in $)

D1 = 0.75 


D2 = 0.75*1.12 = 0.84

D3 = 0.84*1.12 = 0.9408

D4 = 0.9408*1.12 = 1.0537

D5 = 1.0537 *1.12 = 1.1801


D6 = 1.1801*1.02 = 1.2037 


=> P5 = D6 / (Re - G) = 1.2037 / (0.05 - 0.02) = 40.1247 


P0 = 0/75/1.05 + 0.9408/1.05^2 + 1.0537/1.05^3 + (1.1801 + 40.1247)/1.05^5

=> P0 = 34.13 ($) (ANSWER).

answered by: Tulsiram Garg